Grade 12th passMechanics

2) a particle travels half of the time with constt. speed 2m/s in remaining half of the time it travels , 1/4th distance with constt. speed of 4m/s and 3/4th distance with 6m/s find average speed during the complete journey.

sanjeev rana
8 Years agoGrade 12th pass

1 Answer

IHate Ppts
8 Years ago
Divide the question into two parts.
First find out the average speed of 1/4 and 3/4 distance.
Number of different speeds/Average speed=1/V1+1/V2....+1/VN
Applying formula, we get
2/Average speed=1/4+3/6
Average speed=8/3
This average is for second part.
Now, average of both parts is given by the formula
V1+V2/2
We get 2+8/3 whole divided by 2 which is 11/6 m/s.